OpenStudy (anonymous):

PLEAAAAAAAAAAASEEEEEEEEEEEEEE! Explain, in complete sentences, the effect of the difference of squares pattern on the multiplication of radicals. Give an example.

OpenStudy (slaaibak):

The product of a sum and a difference of two values is equal to the difference of their squares. If both of those quantities are square roots of integers, for example p and q such that a = √p and b = √q then performing the expansion squares the radicals leaving only integers. (√p + √q)(√p - √q) = (a + b)(a - b) = a^2 - b^2 = p - q If only one of the quantities is a square root of an integer we get (p + √q)(p - √q) = (a^2 + b)(a^2 - b) = a^4 - b^2 = p^2 - q

OpenStudy (anonymous):

omg youre so wonderful...!!!!!!!!!!